Express the perimeter p of a rectange as a function of the Length L if the length of the rectangle is twice its width.
Answer by stanbon(75887)

Express the perimeter p of a rectangle as a function of the Length L if the length of the rectangle is twice its width.
---
L = 2W
W = L/2
---
Perimeter = 2(length + Width)
P = 2(L + (L/2))
---
P = 2(3/2)L
---
P = 3L
==============
Cheers,
Stan H.